# Magnetic Field Configurations in Binary Neutron Star Mergers I: Post-merger Remnant and Disk

### v1

@@ I Introduction
One important factor influencing the EM counterparts of BNS mergers is the structure and evolution of magnetic fields, since they may shape the post-merger dynamics and drive the launching of outflows. Neutron stars in binary systems have surface magnetic fields of ${\sim}10^{8}-10^{12}\leavevmode\nobreak\ {\rm G}$ [11], which may be amplified during and after the merger up to values of up to ${\sim}10^{16}\leavevmode\nobreak\ {\rm G}$ [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. Such amplification is triggered by the development of magnetohydrodynamic (MHD) instabilities, such as the Kelvin–Helmholtz (KH) instability in the shear layer between the colliding stars [23, 24, 25, 12, 26], and the magnetorotational instability (MRI) in the outer layers of the post-merger remnant and throughout the surrounding disk [27, 17, 14, 28]. Later on, other processes such as magnetic winding can keep amplifying the magnetic field for longer at a lower rate. The highly magnetized post-merger remnant may serve as an engine to launch powerful outflows which can collimate into relativistic jets [16, 29, 30, 17, 31, 32, 33, 34, 35] and thus power sGRBs.
@@ II Methods
POL_UU SR $1.35$ $1.35$ SFHo Pol. Aligned $5$ $28.1$ $8.5\times 10^{-2}$ $2.68\times 10^{-2}$
POL_UD SR $1.35$ $1.35$ SFHo Pol. Antialigned $5$ $31.7$ $9.9\times 10^{-2}$ $2.26\times 10^{-2}$
POL_90d SR $1.35$ $1.35$ SFHo Tilted dipole $90^{\circ}$ $5$ $21.3$ $1.01\times 10^{-1}$ $1.93\times 10^{-2}$
POL_ASYM SR $1.35$ $1.35$ SFHo Pol. displaced $5$ $31.1$ $9.3\times 10^{-2}$ $1.88\times 10^{-2}$
TOR_PP SR $1.35$ $1.35$ SFHo Tor. Aligned $5$ $25.2$ $9.6\times 10^{-2}$ $2.22\times 10^{-2}$
TOR_PM SR $1.35$ $1.35$ SFHo Tor. Antialigned $5$ $28.1$ $7.9\times 10^{-2}$ $3.26\times 10^{-2}$
MIXED SR $1.35$ $1.35$ SFHo Mixed Pol./Tor. $5$ $29.1$ $1.03\times 10^{-1}$ $2.71\times 10^{-2}$
BITANT SR $1.35$ $1.35$ SFHo Pol. (Bitant) $5$ $17.6$ $9.5\times 10^{-2}$ $2.50\times 10^{-2}$
POL_q=1.2 SR $1.47$ $1.22$ SFHo Pol. Aligned $5$ $33.4$ $1.4\times 10^{-1}$ $0.963\times 10^{-2}$
STIFF SR $1.35$ $1.35$ DD2 Pol. Aligned $5$ $21.8$ $1.26\times 10^{-1}$ $2.04\times 10^{-2}$
LOW_B SR $1.35$ $1.35$ SFHo Pol. Aligned $5\times 10^{-7}$ $39.1$ $1.3\times 10^{-1}$ $2.33\times 10^{-2}$
NO_B SR $1.35$ $1.35$ SFHo Hydro 00 $24.2$ $9.9\times 10^{-2}$ $2.56\times 10^{-2}$
POL_UU_HR HR $1.35$ $1.35$ SFHo Pol. Aligned $5$ $13.2$ $8.6\times 10^{-2}$ $2.40\times 10^{-2}$
MIXED_HR HR $1.35$ $1.35$ SFHo Mixed Pol./Tor. $5$ $15.1$ $6.4\times 10^{-2}$ $2.27\times 10^{-2}$
@@
Table 1: List of simulations considered in this study. From left to right we list the assigned name of the simulation, the NS component gravitational masses ($m_{1}$, $m_{2}$), the EOS, the initial configuration of the magnetic field, the initial maximum value of the magnetic field ($\mathcal{B}_{\rm max}^{t=0}$), the duration of the simulation after merger ($t_{\rm end}-t_{\rm mer}$), and the disk ($M_{\rm disk}$) and ejecta ($M_{\rm ej}$) masses at the end of each simulation.
@@
The binary systems’ initial configuration consists of two neutron stars orbiting in a quasi-circular orbit. All but one of our simulations consist of equal-mass NSs with gravitational masses $m_{1}=m_{2}=1.35\leavevmode\nobreak\ M_{\odot}$ at an initial separation of $40$ km; the remaining simulation (POL_q=1.2) evolves an unequal mass binary with NS masses $m_{1}=1.47\leavevmode\nobreak\ M_{\odot}$ and $m_{2}=1.22\leavevmode\nobreak\ M_{\odot}$ (a mass ratio of $q=1.2$) at an initial separation of $45$ km. The initial data was obtained using the Lorene pseudo-spectral code [51], and all but one of our simulations assume the soft SFHo [52] EOS to describe the nuclear properties of the high-density matter; the other simulation (STIFF) assumes the DD2 EOS [53, 54].
@@
On top of the hydrodynamic initial data solution, we superimpose a magnetic field confined within each of the two stars. We consider purely poloidal dipole-like fields, with different orientations of the magnetic moments, purely toroidal fields with varying polarities, and a mixed poloidal/toroidal configuration. For the poloidal cases, we define a purely toroidal vector potential of the form
@@
$\textbf{A}=A_{\rm b}\max(p-p_{\mathrm{cut}},0)^{n_{s}}\\ (x(z^{2}-R^{2}),y(z^{2}-R^{2}),-z(x^{2}+y^{2}-R^{2})),$ (2)
where $R$ is the radius of the stars along the axis of the binary. The normalization constant $A_{\rm b}$ is chosen so that the maximum (conserved) magnetic field, $\mathcal{B}^{i}:=\sqrt{\gamma}B^{i}$, is initially $\mathcal{B}_{\rm max}(t=0)=5\times 10^{15}\leavevmode\nobreak\ {\rm G}$ and it is located at the center of the stars; here, $B^{i}$ are the components of the magnetic field measured by an Eulerian observer and $\gamma:=\det({\gamma_{ij}})$ is the determinant of the spatial metric $\gamma_{ij}$. For comparison, we also evolve a binary without magnetic fields (NO_B) and one with a very low magnetic field of $\mathcal{B}=5\times 10^{8}\leavevmode\nobreak\ {\rm G}$ (LOW_B). In our magnetized cases, the initial magnetic energy, calculated as
@@
Table 1 lists the main setup quantities for all the simulations considered in this study, as well as outcomes such as the final ejecta and disk masses. Note that the values of these latter quantities are affected by the duration of the simulation, especially for the shortest ones.
@@ III Disk and remnant evolution
In this section, we present the results of our analysis of the set of BNS simulations. We first present a qualitative description of the dynamics of the mergers and discuss the main differences among the various setups considered. Then, we quantitatively analyze the evolution of several quantities of interest through the calculation of mass histograms, radial profiles, and the time evolution of averaged values within the remnant and the disk. In what follows, we identify the “remnant” as the region which contains bound material ($-u_{t}<1$) with densities $\rho>10^{13}\leavevmode\nobreak\ {\rm g\leavevmode\nobreak\ cm}^{-3}$, and the “disk” as the region containing bound material with densities in the range $2\times 10^{8}<\rho<10^{13}\leavevmode\nobreak\ {\rm g\leavevmode\nobreak\ cm}^{-3}$.
@@ III.1 Qualitative dynamics
[…]
Binaries with an initial separation of $40$ ($45$) km perform $\approx 3$ ($6$) orbits before merging. We define the merger time as the retarded time for the maximum of the GW $(2,2)$ strain amplitude: $t_{\rm mer}:=t_{\rm max,GW}-R_{\rm GW}/c$, where $R_{\rm GW}\simeq 442\leavevmode\nobreak\ {\rm km}$ is the radius at which GWs are recorded and $c$ is the speed of light. Equal-mass binaries with the SFHo EOS merge at a time of $t_{\rm mer}\simeq 10.3\leavevmode\nobreak\ {\rm ms}$ when using PPM and $t_{\rm mer}\simeq 10.9\leavevmode\nobreak\ {\rm ms}$ when using WENOZ, whereas the STIFF run, with the DD2 EOS, merges at $t_{\rm mer}\simeq 10\leavevmode\nobreak\ {\rm ms}$. On the other hand, simulation POL_q=1.2 merges at $t_{\rm mer}\simeq 15.9\leavevmode\nobreak\ {\rm ms}$, because of the larger initial separation. Depending on the simulation, we continued the evolution for ${\sim}15{-}35\leavevmode\nobreak\ {\rm ms}$ after the merger. All mergers give birth to massive remnants, which survive against gravitational collapse for the entire duration of the runs. We remark that for the masses considered, the binaries with the SFHo EOS are expected to produce a black hole at some time after the merger. Other simulations with a similar setup but performed with a different code have produced a gravitational collapse in a range $\sim 5-25\leavevmode\nobreak\ {\rm ms}$ after the merger (e.g., Refs. [65, 66]). Nevertheless, the collapse time depends on the resolution, as well as on viscous and other dissipative effects, and thus it is not a reliable prediction from simulations [67].
[…]
@@
The first magnetic field structures that exit the remnant are those attached to the material in the spiral arms formed in the late inspiral and early merger (see left panel of Fig. 1). The initially coherent vertical field rapidly spreads out and becomes turbulent. On the contrary, the toroidal field (middle row in Fig. 2) starts from a turbulent state before the merger and slowly becomes more ordered later, especially within the remnant.
@@
The evolution of $\beta^{-1}$ reflects both the amplification and redistribution of the magnetic field. The magnetic field is amplified within the remnant shortly after the merger due to the KH instability (see Paper II). Part of this amplified field is then expelled from the remnant and populates the disk, increasing its magnetic-to-gas pressure ratio there. In the late stages of the simulation, the largest magnetization is present in the outer regions of the disk at moderate angles from the equatorial plane. The bottom row in Fig. 2 also shows that the polar funnel is still dominated by gas pressure by the end of the simulation, indicating that this region is still heavily polluted by baryons. This is consistent with the findings in Ref. [38], who showed that this pollution prevents (or delays) the formation of a relativistic jet. Nevertheless, our treatment neglects neutrino transport, which has been demonstrated to contribute, among other effects, to decrease the baryon pollution in the funnel and then help a jet to form [30, 17, 68].
@@
To compare the effect of the initial binary configuration and magnetic field on the late magnetic structure of the remnant and the disk, we show in Fig. 3 a closer look at the toroidal magnetic field component in the $y=0$ plane for the last snapshot of a subset of our simulations. We have indicated above each panel the time of the last snapshot for each one. Simulations with an initially vertical and aligned dipolar field give rise to a toroidal field that follows the distortion caused by the orbital motion. This produces a counter-clockwise (clockwise) directed field in the regions above (below) the equatorial plane. This is the case for POL_UU, BITANT, POL_q=1.2, and STIFF. On the other hand, simulations with an initially antialigned (POL_UD) dipole-like field or a toroidal field (TOR_PP) develop a toroidal component with the opposite direction above and below the plane compared with the other runs. The simulation with bitant symmetry shows a more coherent structure close to the equatorial plane, which is likely a consequence of the imposed boundary condition, which forces the field to be perfectly symmetric. This suggests that imposing reflective symmetry across the $z=0$ plane decreases the degree of field turbulence that develops close to the plane.
@@
Simulation POL_q=1.2 ran for longer than the other runs and starts to show signs of ordering of the toroidal field in the funnel. Finally, simulation STIFF shows a more rapid ordering of the field within the remnant even at earlier times. This is likely due to two effects: i) as a result of to the stiffness of the EOS, the remnant has a larger volume and thus more space for the formation of coherent structures, and ii) the collision is less violent than with the soft SFHo, and the initial coherent field survives the merger.
@@
Figure 3: Vertical slices ($y=0$) for the last snapshot of a subset of the simulations displaying the component of the magnetic field perpendicular to the plane, $B^{y}$. The BITANT simulation shows a lower development of turbulence near the equatorial plane due to the reflective boundary conditions.
@@
The global structure of the post-merger remnant, disk, and its associated magnetic field can be visualized in Fig. 4, where we show a 3D rendering of the rest-mass density for a late-time simulation snapshot of simulation POL_UD, superposed with the magnetic field lines close to the vertical axes. The red isodensity surfaces correspond to regions where $\rho\geq 10^{13}\leavevmode\nobreak\ {\rm g\leavevmode\nobreak\ cm}^{-3}$, which we defined as the post-merger remnant [69]. Beyond the core, we show in green isodensity surfaces with $10^{12}{\rm g\leavevmode\nobreak\ cm}^{-3}<\rho<10^{13}{\rm g\leavevmode\nobreak\ cm}^{-3}$) and in yellow regions where $2\times 10^{8}{\rm g\leavevmode\nobreak\ cm}^{-3}<\rho<10^{12}{\rm g\leavevmode\nobreak\ cm}^{-3}$; these two regions correspond to the inner and outer disk. The magnetic field structure is visualized using field lines colored by strength, transitioning from yellow (weakest field), through orange, red, and finally purple (strongest field). At this stage of the simulations, the field lines begin to exhibit signs of the characteristic winding caused by the differential rotation of the remnant.
@@
Figure 4: Three-dimensional rendering of rest-mass density and magnetic field structure for the remnant core and surrounding material of simulation POL_UD. The magnetic field lines are colored by strength, transitioning from yellow (weakest field) to purple (strongest field), emphasizing the twisted field structure in the post-merger remnant. The contour volumes show regions of different density, broadly corresponding to the remnant (red, $\rho>10^{13}{\rm g\leavevmode\nobreak\ cm}^{-3}$), inner disk ($10^{12}{\rm g\leavevmode\nobreak\ cm}^{-3}<\rho<10^{13}{\rm g\leavevmode\nobreak\ cm}^{-3}$), and outer disk ($2\times 10^{8}{\rm g\leavevmode\nobreak\ cm}^{-3}<\rho<10^{12}{\rm g\leavevmode\nobreak\ cm}^{-3}$).

### v2

@@ I Introduction
One important factor influencing the EM counterparts of BNS mergers is the structure and evolution of magnetic fields, since they may shape the post-merger dynamics and drive the launching of outflows. Neutron stars in binary systems have surface magnetic fields of ${\sim}10^{8}-10^{12}\leavevmode\nobreak\ {\rm G}$ [11], which may be amplified during and after the merger up to values of up to ${\sim}10^{16}\leavevmode\nobreak\ {\rm G}$ [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. Such amplification is triggered by the development of magnetohydrodynamic (MHD) instabilities, such as the Kelvin–Helmholtz (KH) instability in the shear layer between the colliding stars [23, 24, 25, 12, 26], and the magnetorotational instability (MRI) in the outer layers of the post-merger remnant and throughout the surrounding disk [27, 17, 14, 28]. Subsequently, additional mechanisms such as magnetic winding [29, 30, 20] and the Tayler–Spruit dynamo [31, 32] may further amplify the magnetic field. The latter process, however, also transports angular momentum and may reduce the differential rotation and saturate the toroidal field. The field could rise to the surface of the remnant through the Parker instability [33, 34] and power strong flares, although it remains uncertain whether this occurs unless the field strength becomes excessively large [35]. The highly magnetized post-merger remnant may also serve as an engine to launch powerful outflows which can collimate into relativistic jets [16, 36, 37, 17, 38, 39, 40, 41, 42] and thus power sGRBs.
@@ II Methods
POL-$U$ SR $1.35$ $1.35$ SFHo Pol. Aligned 55 $28.1$ $8.5\times 10^{-2}$ $2.68\times 10^{-2}$
POL-UD SR $1.35$ $1.35$ SFHo Pol. Antialigned 55 $31.7$ $9.9\times 10^{-2}$ $2.26\times 10^{-2}$
POL-90D SR $1.35$ $1.35$ SFHo Tilted dipole ($90^{\circ}$) 55 $21.3$ $1.01\times 10^{-1}$ $1.93\times 10^{-2}$
POL-ASYM SR $1.35$ $1.35$ SFHo Pol. displaced 55 $31.1$ $9.3\times 10^{-2}$ $1.88\times 10^{-2}$
TOR-$P$ SR $1.35$ $1.35$ SFHo Tor. Aligned 55 $25.2$ $9.6\times 10^{-2}$ $2.22\times 10^{-2}$
TOR-PM SR $1.35$ $1.35$ SFHo Tor. Antialigned 55 $28.1$ $7.9\times 10^{-2}$ $3.26\times 10^{-2}$
MIX SR $1.35$ $1.35$ SFHo Mixed Pol./Tor. 55 $29.1$ $1.03\times 10^{-1}$ $2.71\times 10^{-2}$
BITANT SR $1.35$ $1.35$ SFHo Pol. (Bitant) 55 $17.6$ $9.5\times 10^{-2}$ $2.50\times 10^{-2}$
POL-Q12 SR $1.47$ $1.22$ SFHo Pol. Aligned 55 $33.4$ $1.4\times 10^{-1}$ $0.963\times 10^{-2}$
STIFF SR $1.35$ $1.35$ DD2 Pol. Aligned 55 $23.3$ $1.26\times 10^{-1}$ $2.04\times 10^{-2}$
LOWB SR $1.35$ $1.35$ SFHo Pol. Aligned $5\times 10^{-7}$ $39.1$ $1.3\times 10^{-1}$ $2.33\times 10^{-2}$
B0 SR $1.35$ $1.35$ SFHo Hydro 0 $24.2$ $9.9\times 10^{-2}$ $2.56\times 10^{-2}$
HR-POL-$U$ HR $1.35$ $1.35$ SFHo Pol. Aligned 55 $13.2$ $8.6\times 10^{-2}$ $2.40\times 10^{-2}$
HR-MIX HR $1.35$ $1.35$ SFHo Mixed Pol./Tor. 55 $15.1$ $6.4\times 10^{-2}$ $2.27\times 10^{-2}$
@@
Table 1: List of simulations considered in this study. From left to right we list the assigned name of the simulation, the NS component gravitational masses ($m_{1}$, $m_{2}$), the EOS, the initial configuration of the magnetic field, the initial maximum value of the magnetic field ($\mathcal{B}_{\rm max}^{t=0}$), the duration of the simulation after merger ($t_{\rm end}-t_{\rm mer}$), and the disk ($M_{\rm disk}$) and ejecta ($M_{\rm ej}$) masses at the end of each simulation. For the poloidal configurations, by ‘aligned’ (‘antialigned’) we refer to parallel (antiparallel) magnetic moments in both stars, whereas for the toroidal configurations, this term refers to the direction (counterclockwise or clockwise) of the field in both stars. The tilted dipole corresponds to the same topology as in the poloidal runs, but with the magnetic moment rotated $90^{\circ}$ and thus lying on the plane
@@
The binary system’s initial configuration consists of two neutron stars orbiting in a quasi-circular orbit. All but one of our simulations consist of equal-mass NSs with gravitational masses $m_{1}=m_{2}=1.35\leavevmode\nobreak\ M_{\odot}$ at an initial separation of $40$ km; the remaining simulation (POL-Q12) evolves an unequal mass binary with NS masses $m_{1}=1.47\leavevmode\nobreak\ M_{\odot}$ and $m_{2}=1.22\leavevmode\nobreak\ M_{\odot}$ (a mass ratio of $q=1.2$) at an initial separation of $45$ km. The initial data was obtained using the Lorene pseudo-spectral code [58], and all but one of our simulations assume the soft SFHo [59] EOS to describe the nuclear properties of the high-density matter; the other simulation (STIFF) assumes the DD2 EOS [60, 61].
@@
On top of the hydrodynamic initial data solution, we superimpose a magnetic field confined within each of the two stars. We consider purely poloidal dipole-like fields, with different orientations of the magnetic moments, purely toroidal fields with varying polarities, and a mixed poloidal/toroidal configuration. In the latter case, we follow the results in Ref. [62] and assume that $85\%$ of the magnetic energy is in the poloidal field and $15\%$ of it is in the toroidal field. For the poloidal cases, we define a purely toroidal vector potential of the form
@@
$\textbf{A}=A_{\rm b}\max(p-p_{\mathrm{cut}},0)^{n_{s}}\\ \times\left[x(z^{2}-R^{2}),y(z^{2}-R^{2}),-z(x^{2}+y^{2}-R^{2})\right],$ (2)
where $R$ is the radius of the stars along the axis of the binary. The normalization constant $A_{\rm b}$ is chosen so that the maximum (conserved) magnetic field, $\mathcal{B}^{i}:=\sqrt{\gamma}B^{i}$, is initially $\mathcal{B}_{\rm max}(t=0)=5\times 10^{15}\leavevmode\nobreak\ {\rm G}$ and it is located at the center of the stars; here, $B^{i}$ are the components of the magnetic field measured by an Eulerian observer and $\gamma:=\det({\gamma_{ij}})$ is the determinant of the spatial metric $\gamma_{ij}$. For comparison, we also evolve a binary without magnetic fields (B0) and one with a very low magnetic field of $\mathcal{B}=5\times 10^{8}\leavevmode\nobreak\ {\rm G}$ (LOWB). In our magnetized cases, the initial magnetic energy, calculated as
@@
Table 1 lists the main setup quantities for all the simulations considered in this study, as well as outcomes such as the final ejecta (see Paper II) and disk masses. Note that the values of these latter quantities are affected by the duration of the simulation, especially for the shortest ones.
@@ III Disk and remnant evolution
In this section, we present the results of our analysis of the set of BNS simulations. We first present a qualitative description of the dynamics of the mergers and discuss the main differences among the various setups considered. Then, we quantitatively analyze the evolution of several quantities of interest through the calculation of mass histograms, radial profiles, and the time evolution of averaged values within the remnant and the disk. In what follows, we identify the “remnant” as the region which contains bound material ($u_{t}<1$) with densities $\rho>10^{13}\leavevmode\nobreak\ {\rm g\leavevmode\nobreak\ cm}^{-3}$, and the “disk” as the region containing bound material with densities in the range $2\times 10^{8}<\rho<10^{13}\leavevmode\nobreak\ {\rm g\leavevmode\nobreak\ cm}^{-3}$.
@@ III.1 Qualitative dynamics
Figure 1: Slices of various quantities on the equatorial plane for the POL-UU simulation at different times relative to merger time. The panels display the evolution of (top to bottom) the rest-mass density ($\rho$), the vertical component of the magnetic field ($B^{z}$), and the inverse of the plasma $\beta$ parameter. In the bottom two panels, isodensity contours are shown in gray for $\rho=5\times 10^{9},5\times 10^{10},\leavevmode\nobreak\ {\rm and}\leavevmode\nobreak\ 10^{13}\leavevmode\nobreak\ {\rm g\leavevmode\nobreak\ cm}^{-3}$. Each column corresponds to a different time: ($t-t_{\rm mer}=-1.5,0.5,5,15,27.6\leavevmode\nobreak\ {\rm ms}$).
Figure 2: Slices of various quantities on the $xz$ plane for the POL-UU simulation at different times relative to merger time. The panels display the evolution of (top to bottom) the rest-mass density ($\rho$), the magnetic field component perpendicular to the plane ($B^{y}$), and the inverse of the plasma $\beta$ parameter. In the bottom two panels, isodensity contours are shown in gray for $\rho=5\times 10^{9},5\times 10^{10},\leavevmode\nobreak\ {\rm and}\leavevmode\nobreak\ 10^{13}\leavevmode\nobreak\ {\rm g\leavevmode\nobreak\ cm}^{-3}$. Each column corresponds to a different time: ($t-t_{\rm mer}=-1.5,0.5,5,15,27.6\leavevmode\nobreak\ {\rm ms}$).
Binaries with an initial separation of $40$ ($45$) km perform $\approx 3$ (66) orbits before merging. We define the merger time as the retarded time for the maximum of the GW $(2,2)$ strain amplitude: $t_{\rm mer}:=t_{\rm max,GW}-R_{\rm GW}/c$, where $R_{\rm GW}\simeq 442\leavevmode\nobreak\ {\rm km}$ is the radius at which GWs are recorded and $c$ is the speed of light. Equal-mass binaries with the SFHo EOS merge at a time of $t_{\rm mer}\simeq 10.3\leavevmode\nobreak\ {\rm ms}$ when using PPM and $t_{\rm mer}\simeq 11.1\leavevmode\nobreak\ {\rm ms}$ when using WENOZ, whereas the STIFF run, with the DD2 EOS, merges at $t_{\rm mer}\simeq 9\leavevmode\nobreak\ {\rm ms}$. On the other hand, simulation POL-Q12 merges at $t_{\rm mer}\simeq 15.9\leavevmode\nobreak\ {\rm ms}$, because of the larger initial separation. Depending on the simulation, we continued the evolution for ${\sim}15{-}35\leavevmode\nobreak\ {\rm ms}$ after the merger. All mergers give birth to massive remnants, which survive against gravitational collapse for the entire duration of the runs. We remark that for the masses considered, the binaries with the SFHo EOS are expected to produce a black hole at some time after the merger. Other simulations with a similar setup but performed with a different code have produced a gravitational collapse in a range $\sim 5-25\leavevmode\nobreak\ {\rm ms}$ after the merger (e.g., Refs. [73, 74]). Nevertheless, the collapse time depends on the resolution, as well as on viscous and other dissipative effects, and thus it is not a reliable prediction from simulations [75].
Figure 1 shows equatorial slices ($z=0$) at different times for simulation POL-UU, illustrating the time-evolution of representative plasma quantities. From top to bottom, the panels show the rest-mass density ($\rho$), the vertical magnetic field ($B^{z}$), and the magnetic pressure to gas pressure ratio, $\beta^{-1}\equiv p_{B}/p_{\rm gas}$, where $p_{B}=b^{2}/2$ [^2] In some parts of the paper, we refer to this quantity as the ‘magnetization’ because its behavior is very alike that of the $\sigma:=b^{2}/\rho$, but it must not be confused with this latter quantity.. From left to right, each column corresponds to an increasing time relative to the merger time: $t-t_{\rm mer}=-1.5$, $0.5$, 55, $15$, and $27.6$ ms (last snapshot for this particular simulation). In turn, Figure 2 shows equivalent 2D slices but for the $y=0$ plane; in this case, the middle row shows the toroidal component of the magnetic field ($B^{y}$).
@@
The first magnetic field structures that emerge from the remnant are those attached to the material in the spiral arms formed in the late inspiral and early merger (see left panel of Fig. 1). The initially coherent vertical field rapidly spreads out and becomes turbulent. On the contrary, the toroidal field (middle row in Fig. 2) starts from a turbulent state before the merger and slowly becomes more ordered later, especially within the remnant.
@@
The evolution of $\beta^{-1}$ reflects both the amplification and redistribution of the magnetic field. The magnetic field is amplified within the remnant shortly after the merger due to the KH instability (see Paper II). Part of this amplified field is then expelled from the remnant and populates the disk, increasing its magnetic-to-gas pressure ratio there. In the late stages of the simulation, the largest magnetization is present in the outer regions of the disk at moderate angles from the equatorial plane. The bottom row in Fig. 2 also shows that the polar funnel is still dominated by gas pressure by the end of the simulation, indicating that this region is still heavily polluted by baryons. This is consistent with the findings in Ref. [45], who showed that this pollution prevents (or delays) the formation of a relativistic jet. Nevertheless, our treatment neglects neutrino transport, which has been demonstrated to contribute, among other effects, to decreasing the baryon pollution in the funnel and then help a jet to form [37, 17, 76].
@@
To compare the effect of the initial binary configuration and magnetic field on the late magnetic structure of the remnant and the disk, we show in Fig. 3 a closer look at the toroidal magnetic field component in the $y=0$ plane for the last snapshot of a subset of our simulations. We have indicated above each panel the time of the last snapshot for each one. Simulations with an initially vertical and aligned dipolar field give rise to a toroidal field that follows the natural twisting caused by the rotation of the remnant, namely, this produces a counter-clockwise (clockwise) directed field in the regions above (below) the equatorial plane. This is the case for POL-UU, BITANT, POL-Q12, and STIFF, which share the same initial field configuration. On the other hand, simulations with a different initial configuration, such as initially antialigned (POL-UD) dipole-like fields or toroidal fields (TOR-PP), develop a toroidal component with the opposite direction above and below the plane compared with the aforementioned runs. The POL-UD configuration is of particular interest, as antialigned fields do not prescribe a preferred direction for the post-merger toroidal component. Accordingly, the topology we see in our simulation may not constitute a generic outcome of this setup, but rather reflect a stochastic realization influenced by subtle asymmetries in the initial conditions.
Figure 4 shows a magnified comparison between the toroidal magnetic field in the POL-UU and BITANT runs. The simulation with bitant symmetry exhibits a more coherent structure near the equatorial plane, which is likely a consequence of the imposed boundary condition that forces the field to be perfectly symmetric. This suggests that imposing reflective symmetry across the $z=0$ plane decreases the degree of field turbulence that develops close to the plane.
@@
Simulation POL-Q12 ran for longer than the other runs and starts to show signs of ordering of the toroidal field in the funnel. Finally, simulation STIFF shows a more rapid ordering of the field within the remnant even at earlier times. This is likely due to two effects: i) as a result of to the stiffness of the EOS, the remnant has a larger volume and thus more space for the formation of coherent structures, and ii) the collision is less violent than with the soft SFHo, and the initial coherent field survives the merger.
@@
Figure 3: Vertical slices ($y=0$) for the last snapshot of a subset of the simulations displaying the component of the magnetic field perpendicular to the plane, $B^{y}$.
Figure 4: Vertical slices ($y=0$) at $t=15\leavevmode\nobreak\ {\rm ms}$ for POL-UU and BITANT, displaying the component of the magnetic field perpendicular to the plane, $B^{y}$. The BITANT simulation shows a lower development of turbulence near the equatorial plane due to the reflective boundary conditions.
@@
The global structure of the post-merger remnant, disk, and its associated magnetic field can be visualized in Fig. 5, where we show a 3D rendering of the rest-mass density for a late-time simulation snapshot of simulation POL-UD, superposed with the magnetic field lines close to the vertical axes. The red isodensity surfaces correspond to regions where $\rho\geq 10^{13}\leavevmode\nobreak\ {\rm g\leavevmode\nobreak\ cm}^{-3}$, which we defined as the post-merger remnant [77]. Beyond the core, we show in green isodensity surfaces with $10^{12}{\rm g\leavevmode\nobreak\ cm}^{-3}<\rho<10^{13}{\rm g\leavevmode\nobreak\ cm}^{-3}$) and in yellow regions where $2\times 10^{8}{\rm g\leavevmode\nobreak\ cm}^{-3}<\rho<10^{12}{\rm g\leavevmode\nobreak\ cm}^{-3}$; these two regions correspond to the inner and outer disk. The magnetic field structure is visualized using field lines colored by strength, transitioning from yellow (weakest field), through orange, red, and finally purple (strongest field). At this stage of the simulations, the field lines begin to exhibit signs of the characteristic winding caused by the differential rotation of the remnant.
@@
Figure 5: Three-dimensional rendering of rest-mass density and magnetic field structure for the remnant core and surrounding material of simulation POL-UD. The magnetic field lines are colored by strength, transitioning from yellow (weakest field) to purple (strongest field), emphasizing the twisted field structure in the post-merger remnant. The contour volumes show regions of different density, broadly corresponding to the remnant (red, $\rho>10^{13}{\rm g\leavevmode\nobreak\ cm}^{-3}$), inner disk ($10^{12}{\rm g\leavevmode\nobreak\ cm}^{-3}<\rho<10^{13}{\rm g\leavevmode\nobreak\ cm}^{-3}$), and outer disk ($2\times 10^{8}{\rm g\leavevmode\nobreak\ cm}^{-3}<\rho<10^{12}{\rm g\leavevmode\nobreak\ cm}^{-3}$).